# Adaptive Incentive Design with Learning Agents

### v2

@@ 4.1 Atomic Aggregative Games
First, we show that if $M$ is invertible, then ${x^{\ast}}(p)=-M^{-1}p$ for any $p\in\mathbb{R}^{|\mathcal{I}|}$. To see this, observe that the cost function $c_{i}(x_{i},x_{-i},p)$ is strongly convex in $x_{i}$ and the strategy space $X_{i}$ is unconstrained. Therefore, it must hold that $\nabla_{x_{i}}c_{i}(x^{\ast}(p),p)=0$, for every $i\in\mathcal{I}$. Consequently, using (28), it must hold that
Stacking (30) in a vector form yields $Mx^{\ast}(p)=-p$.
Next, we show that $P^{\dagger}$ is a singleton set. From (7a), the externality in this setup is expressed as:
$e_{i}(x)=\nabla h_{i}(x_{i})-q_{i}x_{i}-\alpha\sum_{j=1}^{n}A_{ij}x_{j},~{}\forall i\in\mathcal{I},~{}\forall x.$ (31)
@@
$e_{i}(x^{\ast}(p))=\nabla h_{i}(x_{i}^{\ast}(p))+p_{i}.$ (32)
Consequently, using (LABEL:eq:_PDagger), we obtain
@@
Since $h_{i}$ is strictly convex, from Assumption 4.1, there exists a unique $y^{\dagger}$ such that $\nabla h_{i}(y_{i}^{\dagger})=0$, for every $i\in I$. Therefore, for every $p^{\dagger}\in P^{\dagger}$ it should hold that $x^{\ast}(p^{\dagger})=y^{\dagger}$. Using the fact that $x^{\ast}(p)=-M^{-1}p$, we obtain the unique $p^{\dagger}=-My^{\dagger}$. ∎
Next, we provide sufficient conditions to ensure the convergence of ($x$-update)-($p$-update).
Proposition 4.2.
@@
Consider the aggregative game $G$ with invertible $M$. Suppose that Assumptions 3.1, 3.2, and 4.1 are satisfied, and at least one of the following conditions hold:
  • (A1)
    The matrix $M$ has non-negative entries, $M^{-1}$ has strictly-negative off-diagonal entries. Furthermore, there exists [^4] Since $h_{i}$ is strictly convex then $\nabla h_{i}(\cdot)$ is strictly increasing and can have only one zero-crossing, which ensures that $y^{\dagger}$ is unique. some $y^{\dagger}\in\mathbb{R}^{|\mathcal{I}|}$ such that $\nabla h_{i}(y^{\dagger}_{i})=0$ for every $i\in\mathcal{I}$, and all entries of $y^{\dagger}$ have the same sign;
  • (A2)
    $M$ is a symmetric positive definite matrix.
Additionally, $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$. Then there exist positive scalars $\bar{r},\bar{\alpha},\bar{\beta},\bar{\gamma}$ such that for every $p_{0}\in\mathcal{B}_{\bar{r}}(p^{\dagger})$, $x_{0}\in\mathcal{B}_{\bar{r}}(x^{\ast}(p_{0}))$, with step-sizes that follow (18) must satisfy $\lim_{k\to\infty}(x_{k},p_{k})=({x}^{\dagger},p^{\dagger})$.
Proof.
The proof follows by verifying that the conditions $(A1)$ and $(A2)$ ensure that the conditions $(C1)$ and $(C2)$ in Proposition 3.3 are satisfied, respectively.
To begin, we show that condition $(A1)$ ensures that condition $(C1)$ is satisfied in Proposition 3.3. Recall, from (32), that $e_{i}(x^{\ast}(p))=\nabla h_{i}(x_{i}^{\ast}(p))+p_{i}$. First, we show that $i,j\in\mathcal{I}$ such that $i\neq j,$ $\frac{\partial e_{i}(x^{\ast}(p))}{\partial p_{j}}>0$. Indeed,
@@
where the inequality holds because $h_{i}$ is strictly convex and $(M^{-1})_{ij}<0$.
Next, we show that if $e_{i}(x^{\ast}(0))\geqslant 0$ for every $i\in\mathcal{I}$ then there exists $p\in\mathbb{R}_{+}^{|\mathcal{I}|}$ such that $e_{i}(x^{\ast}(p))-p_{i}\leqslant 0$ for every $i\in\mathcal{I}$ (cf. (C1)-(i) in Proposition 3.3). Using (32), it is sufficient to show that
[…]
$\nabla h_{i}(x^{\ast}(\bar{p}))\leqslant 0\quad\forall\ i\in\mathcal{I}.$
By $(A1)$, we know that $\nabla h_{i}(y^{\dagger}_{i})=0$ and $y^{\dagger}_{i}y^{\dagger}_{j}\geqslant 0$ for every $i,j\in\mathcal{I}$. We claim that $y^{\dagger}_{i}\nabla h_{i}(0)\leqslant 0$ for every $i\in\mathcal{I}$. Indeed, for every $i\in\mathcal{I}$,
$0\leqslant(\nabla h_{i}(y^{\dagger}_{i})-\nabla h_{i}(x^{\ast}_{i}(0)))(y^{\dagger}_{i}-x_{i}^{\ast}(0))=-y^{\dagger}_{i}\nabla h_{i}(0),$
where the first inequality is due to the strict convexity of $h_{i}$. Suppose $\nabla h_{i}(0)\geqslant 0$ , for every $i\in\mathcal{I}$, then $-y^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$. Next, define $\bar{p}=-(1+\epsilon)My^{\dagger}$, for some $\epsilon>0$, then $\bar{p}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ as $y^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ and $M$ has all non-negative entries (cf. $(A1)$). This ensures that $x^{\ast}(\bar{p})=(1+\epsilon)y^{\dagger}$. We claim that $\nabla h_{i}(x^{\ast}_{i}(\bar{p}))<0$ for every $i\in\mathcal{I}$. Indeed, for every $i\in\mathcal{I}$,
00 0 $<(\nabla h_{i}(x_{i}^{\ast}(\bar{p}))-\nabla h_{i}(y^{\dagger}_{i}))(x_{i}^{\ast}(\bar{p})-y^{\dagger}_{i})$
$=\nabla h_{i}(x_{i}^{\ast}(\bar{p}))\epsilon y^{\dagger}_{i}.$
The claim follows because $y^{\dagger}_{i}\leqslant 0$ and $\epsilon>0$. This shows that the implication in (34) holds. Similarly, it can be show that if $e_{i}(x^{\ast}(0))\leqslant 0$ for every $i\in\mathcal{I}$ then there exists $p\in\mathbb{R}_{-}^{|\mathcal{I}|}$ such that $e_{i}(x^{\ast}(p))-p_{i}\geqslant 0$ for every $i\in\mathcal{I}$ (cf. (C1)-(ii) in Proposition 3.3).
Next, we verify that condition $(A2)$ ensures that condition $(C2)$ in Proposition 3.3 is satisfied. Consider the function $V(p)=(p-p^{\dagger})^{\top}M^{-\top}(p-p^{\dagger})$. Note that $V(p^{\dagger})=0$ and $V(p)>0$ for all $p\neq p^{\dagger}$. Next, we show that $\nabla V(p)^{\top}(e(x^{\ast}(p))-p)\leqslant 0.$ Indeed,
$=-2(x^{\ast}(p)-x^{\ast}(p^{\dagger}))\nabla h(x^{\ast}(p))$
$\leqslant 0$
where the last equality holds due to the strict convexity of $h_{i}$ for every $i\in\mathcal{I}$. ∎
Remark 4.3.
There exists instances of aggregative game such that only one of the conditions $(A1)$ or $(A2)$ is satisfied in Proposition 4.2. For instance consider the matrices
$M_{1}=\begin{bmatrix}1&0.1\\ 1&1\end{bmatrix},M_{2}=\begin{bmatrix}1&-0.1\\ -0.1&1\end{bmatrix}.$
Note that $M_{1}$ satisfies $(A1)$ but not $(A2)$ and the matrix $M_{2}$ satisfies $(A2)$ but not $(A1)$.
4.2 Non-atomic Traffic Routing on General Networks
Consider a routing game $\tilde{G}$ characterizing the interaction of strategic travelers over a graph $\tilde{\mathcal{G}}=(\tilde{\mathcal{E}},\tilde{\mathcal{N}})$ where $\tilde{\mathcal{N}}$ is the set of nodes and $\tilde{\mathcal{E}}$ denotes the set of edges. Let $\tilde{\mathcal{I}}$ be the set of origin-destination (o-d) pairs. Each o-d pair $i\in\tilde{\mathcal{I}}$ is connected by a set of routes [^5] A route is a set of contiguous edges. represented by $\tilde{R}_{i}$. Let $\tilde{R}=\cup_{i\in\tilde{\mathcal{I}}}\tilde{R}_{i}$ denote the set of all routes on the network.
An infinitesimal traveler on the network is associated with an o-d pair and chooses a route to commute between the o-d pair. Let the total population of travelers associated with any o-d pair $i\in\tilde{\mathcal{I}}$ be denote by $\tilde{M}_{i}$. Let $\tilde{x}_{i}^{j}$ be the amount of travelers taking route $j\in\tilde{R}_{i}$ to commute between o-d pair $i\in\tilde{\mathcal{I}}$ and $\tilde{x}=(\tilde{x}_{i}^{j})_{j\in\tilde{R}_{i},i\in\tilde{\mathcal{I}}}$ is a vector which contains, as its entries, the route flow of all population on different routes. Naturally, it holds that $\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}=\tilde{M}_{i}$ for every $i\in\tilde{\mathcal{I}}$. Any route flow $\tilde{x}$ induces a flow on the edges of the network denoted by $\tilde{w}$ such that $\tilde{w}_{a}=\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}\mathbbm{1}(a\in j)$ for every $a\in\tilde{\mathcal{E}}$. We denote the set of feasible route flows by $\tilde{X}$ and the set of feasible edge flows by $\tilde{W}=\{(\tilde{w}_{a})_{a\in\tilde{\mathcal{E}}}:\exists\tilde{x}\in\tilde{X},\tilde{w}_{a}=\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}\}$.
For any o-d pair $i\in\tilde{\mathcal{I}}$ and route flow $\tilde{x}\in\mathbb{R}^{|\tilde{R}|}$, the cost experienced by travelers using route $j\in\tilde{R}_{i}$ is ~$\tilde{\ell}_{i}^{j}(\tilde{x})=\sum_{a\in\tilde{\mathcal{E}}}l_{a}(\tilde{w}_{a})\mathbbm{1}(a\in j)$, where $l_{a}(\cdot)$ is the edge latency function that depends on the edge flows. For every edge $a\in\tilde{\mathcal{E}}$, we assume that the edge latency function $l_{a}(\cdot)$ is convex and strictly increasing. This property of edge latency function captures the congestion effect on the transportation network [Beckmann et al., 1956, Roughgarden, 2010].
A social planner designs incentives by setting tolls on the edges of the network in the form of edge tolls [^6] If we directly use the setup of non-atomic games presented in Section 2.1.2 we would require the social planner to use route-based tolls rather than edge-based tolls. Our approach of using edge-based tolls is rooted in practical consideration with implementation of tolls., denoted by $\tilde{p}=(\tilde{p}_{a})_{a\in\tilde{\mathcal{E}}}$. Every edge toll vector induces a unique route toll vector $\tilde{P}$. That is, for any o-d pair $i\in\tilde{\mathcal{I}}$ the toll on route $j\in\tilde{R}_{i}$ is
$\tilde{P}_{i}^{j}=\sum_{a\in\tilde{\mathcal{E}}:a\in j}\tilde{p}_{a}.$ (35)
Consequently, the total cost experienced by travelers on o-d pair $i\in\tilde{\mathcal{I}}$ who choose route $j\in\tilde{R}_{i}$ is $\tilde{c}_{i}^{j}(\tilde{x},\tilde{P})=\tilde{\ell}_{i}^{j}(\tilde{x})+\tilde{P}_{i}^{j}$. Let $\tilde{x}^{\ast}(\tilde{P})$ denote a Nash equilibrium corresponding to route tolls $\tilde{P}$. Owing to (35), with slight abuse of notation, we shall frequently use $\tilde{x}^{\ast}(\tilde{P})$ and $\tilde{x}^{\ast}(\tilde{p})$ interchangeably. Typically, the equilibrium route flows can be non-unique but the corresponding edge flows $\tilde{w}^{\ast}(\tilde{p})$ are unique. Furthermore, the function $\tilde{p}\mapsto\tilde{w}^{\ast}(\tilde{p})$ is a continuous function [Yang and Huang, 2005].
The social planner’s objective is to design tolls that ensure that the resulting equilibrium minimizes the overall travel time incurred by travelers on the network, characterized as the minimizer of
Using the description of cost of travelers, the externality, based on (7b), caused by travelers from o-d pair $i\in\tilde{\mathcal{I}}$ using the route $j\in\tilde{R}_{i}$ is
where $(a)$ is due to Lemma B.2 in Appendix B.
Define
$\tilde{\textbf{P}}^{\dagger}=\{(\tilde{p}^{\dagger}_{a})_{a\in\tilde{\mathcal{E}}}:\tilde{p}^{\dagger}_{a}=\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})),\forall\ a\in\tilde{\mathcal{E}}\big{\}}.$
Proposition 4.4.
Proof.
@@ 4.2 Non-atomic Traffic Routing on General Networks
First, we show that $\tilde{\textbf{P}}^{\dagger}$ is non-empty. This can be shown analogous to the proof of existence in Proposition 3.1 by using Schauder fixed point theorem and continuity of function $\tilde{w}^{\ast}(\cdot)$. We omit the details of this proof for the sake of brevity.
@@
Next, we show that any $p^{\dagger}\in\tilde{\textbf{P}}^{\dagger}$ aligns the Nash equilibrium with social optimality, i.e. $\tilde{w}(p^{\dagger})=w^{\dagger}$. For any $p^{\dagger}\in\tilde{\textbf{P}}^{\dagger}$, we have $\tilde{p}^{\dagger}_{a}=\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger}))$ for every $a\in\tilde{\mathcal{E}}$. This implies,
$\frac{\partial}{\partial\tilde{w}_{a}}(\tilde{w}_{a}(\tilde{p}^{\dagger})l_{a}(\tilde{w}_{a}(\tilde{p}^{\dagger})))=l_{a}(\tilde{w}_{a}(\tilde{p}^{\dagger}))+\tilde{p}^{\dagger}_{a},\quad\forall\ a\in\tilde{\mathcal{E}}.$ (38)
@@
$\sum_{a\in\tilde{\mathcal{E}}}\left(l_{a}(\tilde{w}_{a}(\tilde{p})+\tilde{p}_{a})(\tilde{w}_{a}-\tilde{w}_{a}(\tilde{p}))\right)\geqslant 0\quad\forall\ \tilde{w}\in\tilde{W}.$ (40)
Combining (38) and (40), we conclude that
$\sum_{a\in\tilde{\mathcal{E}}}\frac{\partial}{\partial\tilde{w}_{a}}(\tilde{w}_{a}(\tilde{p}^{\dagger})l_{a}(\tilde{w}_{a}(\tilde{p}^{\dagger})))(\tilde{w}_{a}-\tilde{w}_{a}^{\ast}(\tilde{p}^{\dagger}))\geqslant 0\quad\forall\ \tilde{w}\in\tilde{W}.$ (41)
Further, from the first order conditions of optimality for social cost function, we know that $\tilde{x}^{\dagger}$ is socially optimal if any only if
@@
Using Lemma B.3, we can equivalently write (42) in terms of edge flows as follows
@@
where $w^{\dagger}$ is the edge flow corresponding to the route flow $x^{\dagger}$. Comparing (41) with (43), we note that $\tilde{w}^{\ast}(p^{\dagger})$ is the minimizer of social cost function $\tilde{\Phi}$. This implies $\tilde{w}^{\ast}(\tilde{p}^{\dagger})=\tilde{w}^{\dagger}$.
The proof that $\tilde{\textbf{P}}^{\dagger}$ is singleton follows by contradiction which is analogous to that in Proposition 3.1. We omit the details for the sake of brevity.
Next, we show that the updates ($\tilde{x}$-update)-($\tilde{p}$-update) converge to the fixed points discussed in previous section.
Proposition 4.5.
Suppose that for every $i\in\tilde{I},j\in\tilde{R}_{i}$
$\tilde{c}_{i}^{j}(\tilde{x}^{\ast}(\tilde{p}^{\dagger}))\leqslant\tilde{c}_{i}^{j^{\prime}}(\tilde{x}^{\ast}(\tilde{p}^{\dagger}))\ \forall\ j^{\prime}\in\tilde{R}_{i}\implies\tilde{x}_{i,j}^{\ast}(\tilde{p}^{\dagger})>0.$ (44)
Additionally, $\sup_{k\in\mathbb{N}}(\|\tilde{x}_{k}\|+\|\tilde{p}_{k}\|)<+\infty$. Then there exist positive scalars $\bar{r},\bar{\alpha},\bar{\beta},\bar{\gamma}$ such that for every $\tilde{p}_{0}\in\mathcal{B}_{\bar{r}}(\tilde{p}^{\dagger})$, $\tilde{x}_{0}\in\mathcal{B}_{\bar{r}}(\tilde{x}^{\ast}(\tilde{p}_{0}))$, with step-sizes that follow (18) must satisfy $\lim_{k\to\infty}(\tilde{x}_{k},\tilde{p}_{k})=(\tilde{x}^{\dagger},\tilde{p}^{\dagger})$.
Proof.
The proof follows by verifying the second convergence condition in Proposition 3.3. Particularly, we show that $V(\tilde{p})=(\tilde{p}-\tilde{p}^{\dagger})^{\top}\Delta(\tilde{p}-\tilde{p}^{\dagger})$ serves as candidate Lyapunov function which satisfies the conditions stated in Proposition 3.3-(C2) if we set $\Delta\in\mathbb{R}^{|\tilde{\mathcal{E}}|\times|\tilde{\mathcal{E}}|}$ to be a diagonal matrix such that
$\Delta_{a,a}=(\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger}))+\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})\nabla^{2}l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})))^{-1}.$
Due to the strict monotonicity and convexity of $l_{a}(\cdot)$, it follows that $\Delta_{a,a}>0$ for every $a\in\tilde{\mathcal{E}}$. Towards this goal, we show that there exists a positive scalar $r$ such that for any $\tilde{p}\in\mathcal{B}_{r}(\tilde{p}^{\dagger})$, it holds that
@@
$=-2V(\tilde{p})+2\sum_{a\in\tilde{\mathcal{E}}}\Delta_{a,a}(\tilde{p}_{a}-\tilde{p}^{\dagger}_{a})\left(\phi_{a}(\tilde{p})-\phi_{a}(\tilde{p}^{\dagger})\right)$
@@
where for every $a\in\tilde{\mathcal{E}}$, $\phi_{a}(\tilde{p}):=\tilde{w}^{\ast}_{a}(\tilde{p})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}))$. Thus it suffices to show that
@@
$\sum_{a\in\tilde{\mathcal{E}}}\Delta_{a,a}(\tilde{p}_{a}-\tilde{p}^{\dagger}_{a})\left(\phi_{a}(\tilde{p})-\phi_{a}(\tilde{p}^{\dagger})\right)\leqslant 0.$ (46)

### v3

@@ B.3 Additional Results
The proof that $x^{\ast}(p)=-M^{-1}p$ follows exactly as in Proposition 4.2. Next, we show that $P^{\dagger}$ is a singleton. Using (8a), (29), and (44), the externality is given by
$e_{i}(x)=\nabla h_{i}(x_{i})-q_{i}x_{i}-\alpha\sum_{j\in\mathcal{I}}A_{ij}x_{j},\quad\forall i\in\mathcal{I}.$ (45)
@@
$e_{i}(x^{\ast}(p))=\nabla h_{i}(x_{i}^{\ast}(p))+p_{i},\quad\forall i\in\mathcal{I}.$ (46)
Consequently, using (LABEL:eq:_PDagger), we have
@@
Since $h_{i}$ is strictly convex, it follows from Assumption B.1 that there exists a unique $y^{\dagger}$ such that $\nabla h_{i}(y_{i}^{\dagger})=0$ for every $i\in\mathcal{I}$. Therefore, for every $p^{\dagger}\in P^{\dagger}$, it must hold that $x^{\ast}(p^{\dagger})=y^{\dagger}$. Since $x^{\ast}(p)=-M^{-1}p$, it follows that $p^{\dagger}=-My^{\dagger}$, establishing the uniqueness of $P^{\dagger}$. ∎
Next, we provide sufficient conditions to ensure the convergence of ($x$-update)-($p$-update) to the fixed points. In particular, we present two sets of conditions: the first set establishes global convergence guarantees, while the second set ensures local convergence guarantees.
Proposition B.2.
@@
Consider the updates ($x$-update)-($p$-update) associated with the aggregative game $G$. Suppose that Assumptions 3.1, 3.2, and B.1 hold, and that $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$. Additionally,
  • (i)
    If $M:=Q+\alpha A$ is symmetric positive definite, then the discrete-time updates ($x$-update) and ($p$-update) globally converge to the fixed point $({x}^{\dagger},p^{\dagger})$ in the sense of Definition 3.2.
  • (ii)
    If $M:=Q+\alpha A$ is invertible with non-negative entries, $M^{-1}$ has strictly negative off-diagonal entries, and there exists a vector $y^{\dagger}\in\mathbb{R}^{|\mathcal{I}|}_{-}$ such that $\nabla h_{i}(y^{\dagger}_{i})=0$ for every $i\in\mathcal{I}$  [^9] A similar statement can be obtained for the case when $y^{\dagger}\in\mathbb{R}^{|\mathcal{I}|}_{+}$, but we omit it for brevity. then the discrete-time updates ($x$-update) and ($p$-update) locally converge to the fixed point $({x}^{\dagger},p^{\dagger})$ in the sense of Definition 3.2.
Propositions B.2-(i) and 4.2 are related but differ in two key aspects. First, the social cost function in Proposition 4.2 is a special case of the more general function in (44). Second, Proposition B.2-(i) directly assumes boundedness of iterates, $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$, whereas Proposition 4.2 instead relies on the global convergence of the limiting dynamical system associated with strategy updates (cf. (31)). The simpler social cost function (30) in Proposition 4.2 allows us to use stability results from two-timescale approximation theory [Lakshminarayanan and Bhatnagar, 2017, Theorem 10] to establish boundedness. Extending this approach to Proposition B.2 would require imposing global convergence of suitably defined limiting dynamical systems (cf. (42)-(43)). To maintain clarity, we impose $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$ directly in Proposition B.2.
The conditions imposed on the matrix $M$ in Proposition B.2-(i) and (ii) are not directly comparable; neither necessarily implies the other 101010 For example, consider aggregative games with parameters $(Q_{1},A_{1})$ and $(Q_{2},A_{2})$ such that $M_{1}=Q_{1}+\alpha A_{1}$, $M_{2}=Q_{2}+\alpha A_{2}$ and $M_{1}=\begin{bmatrix}1&0.1\\ 1&1\end{bmatrix},\quad M_{2}=\begin{bmatrix}1&-0.1\\ -0.1&1\end{bmatrix}.$ The matrix $M_{1}$ satisfies the conditions in Proposition B.2-(ii) but does not satisfy the conditions in Proposition B.2-(i). On the other hand, the matrix $M_{2}$ satisfies the conditions in Proposition B.2-(i) but does not satisfy the conditions in Proposition B.2-(ii)..
Proof of Proposition B.2: We prove Proposition B.2(i)-(ii) in order.
  • (a)
    Proposition B.2-(i) follows by verifying the requirements (R1’)-(R2) of Proposition 3.3. We only need to verify (R1’) as (R2) is satisfied due to the assumption that $\sup_{k}\|x_{k}\|+\|p_{k}\|<\infty$. We define a Lyapunov function candidate $V(p)=(p-p^{\dagger})^{\top}M^{-\top}(p-p^{\dagger})$ for the dynamical system (20). Note that $V(p^{\dagger})=0$ and $V(p)>0$ for all $p\neq p^{\dagger}$. Next, we show that $\nabla V(p)^{\top}(e(x^{\ast}(p))-p)<0$ for every $p\neq p^{\dagger}$. Indeed,
    $=-2(x^{\ast}(p)-x^{\ast}(p^{\dagger}))\nabla h(x^{\ast}(p))$
[…]
    $<0,\quad\forall\ p\neq p^{\dagger},$
    where the last equality follows from the strict convexity of $h_{i}$ for each $i\in\mathcal{I}$, completing the proof.
  • (b)
    Proposition B.2-(ii) follows by verifying conditions (R1) and (R2) of Proposition 3.3. Given that $\sup_{k}\|x_{k}\|+\|p_{k}\|<\infty$, it suffices to verify (R1). This follows since condition (C1) in Lemma D.1 holds under Proposition 4.2-(ii) and Assumption B.1.
    First, we show that for $i,j\in\mathcal{I}$ with $i\neq j$, it holds that $\frac{\partial e_{i}(x^{\ast}(p))}{\partial p_{j}}>0$. Indeed,
@@
    where the inequality follows from the strict convexity of $h_{i}$ and the fact that $(M^{-1})_{ij}<0$. Second, we show that condition (C1)-(i) in Lemma D.1 holds. First, we establish that $e_{i}(x^{\ast}(0))\geqslant 0$ for every $i\in\mathcal{I}$. From (46), we note that $e_{i}(x^{\ast}(0))=\nabla h_{i}(0)$ for every $i\in\mathcal{I}$. Therefore, it suffices to show that $\nabla h_{i}(0)\geqslant 0$ for all $i\in\mathcal{I}$. By Assumption B.1, $\nabla h_{i}(\cdot)$ is strictly increasing, and for each $i\in\mathcal{I}$, there exists a unique $y_{i}^{\dagger}\leqslant 0$ such that $\nabla h_{i}(y_{i}^{\dagger})=0$. This implies that $\nabla h_{i}(0)\geqslant 0$, for every $i\in\mathcal{I}$. Next, we verify that $p^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$. From Proposition B.1, $p^{\dagger}=-My^{\dagger}$. Since $M$ has non-negative entries and $y^{\dagger}\in\mathbb{R}_{-}^{|\mathcal{I}|}$, it follows that $p^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$. Finally, we show the other condition in (C1)-(i) in Lemma D.1, which requires that for any $p\in\mathbb{R}_{+}^{|\mathcal{I}|}$, there exists $p^{\prime}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ such that for every $i\in\mathcal{I}$, $p^{\prime}_{i}>p_{i}$ and $e_{i}(x^{\ast}(p^{\prime}))-p_{i}^{\prime}\leqslant 0$, for all $i\in\mathcal{I}$. To show this, we define $p^{\epsilon}=-(1+\epsilon)My^{\dagger}$ for every $\epsilon>0$. Note that $p^{\epsilon}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ and for any $p\in\mathbb{R}^{|\mathcal{I}|}_{+}$, we can select $\epsilon>0$ such that $p^{\epsilon}_{i}>p_{i}$ for every $i\in\mathcal{I}$. Therefore, we show that for every $\epsilon>0$,
    $e_{i}(x^{\ast}(p^{\epsilon}))-p^{\epsilon}\leqslant 0,\quad\forall i\in\mathcal{I}.$ (48)
    From (46), we note that $e_{i}(x^{\ast}(p^{\epsilon}))-p^{\epsilon}=\nabla h_{i}(x^{\ast}_{i}(p^{\epsilon})),$ for every $i\in\mathcal{I}.$ Therefore, to show (48), it is sufficient to show that
    $\nabla h_{i}(x^{\ast}_{i}(p^{\epsilon}))\leqslant 0,\quad\forall i\in\mathcal{I},\epsilon>0.$ (49)
    Indeed, for every $i\in\mathcal{I}$ and $\epsilon>0$,
    00 0 $<(\nabla h_{i}(x_{i}^{\ast}({p}^{\epsilon}))-\nabla h_{i}(y^{\dagger}_{i}))(x_{i}^{\ast}(p^{\epsilon})-y^{\dagger}_{i})$
    $=\nabla h_{i}(x_{i}^{\ast}(p^{\epsilon}))\epsilon y^{\dagger}_{i},$
    where we note that $x^{\ast}(p^{\epsilon})=(1+\epsilon)y^{\dagger}$. To conclude, (49) follows because $y^{\dagger}_{i}\leqslant 0$ and $\epsilon>0$.
Appendix C Proofs of Results in Section 4.2
C.1 Proof of Proposition 4.3.
@@
First, we show that $\tilde{\textbf{P}}^{\dagger}$ is non-empty. This can be shown analogously to the proof of existence in Proposition 3.1 by using the Schauder fixed-point theorem and the continuity of the function $\tilde{w}^{\ast}(\cdot)$. We omit the details of this proof for the sake of brevity.
@@
Next, we show that any $p^{\dagger}\in\tilde{\textbf{P}}^{\dagger}$ aligns the Nash equilibrium with social optimality, i.e. $\tilde{w}(p^{\dagger})=w^{\dagger}$. For any $p^{\dagger}\in\tilde{\textbf{P}}^{\dagger}$, we have $\tilde{p}^{\dagger}_{a}=\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger}))$ for every $a\in\tilde{\mathcal{E}}$. This implies, for every $a\in\tilde{\mathcal{E}},$
$\frac{\partial}{\partial\tilde{w}_{a}}\left(\tilde{w}_{a}(\tilde{p}^{\dagger})l_{a}(\tilde{w}_{a}(\tilde{p}^{\dagger}))\right)=l_{a}(\tilde{w}_{a}(\tilde{p}^{\dagger}))+\tilde{p}^{\dagger}_{a}.$ (50)
@@
$\sum_{a\in\tilde{\mathcal{E}}}\Big{(}l_{a}\big{(}\tilde{w}_{a}(\tilde{p})+\tilde{p}_{a}\big{)}\big{(}\tilde{w}_{a}-\tilde{w}_{a}(\tilde{p})\big{)}\Big{)}\geqslant 0,~{}\forall\ \tilde{w}\in\tilde{W}.$ (52)
Combining (50) and (52), we conclude that for every $\tilde{w}\in\tilde{W},$
$\sum_{a\in\tilde{\mathcal{E}}}\frac{\partial}{\partial\tilde{w}_{a}}(\tilde{w}_{a}(\tilde{p}^{\dagger})l_{a}(\tilde{w}_{a}(\tilde{p}^{\dagger})))(\tilde{w}_{a}-\tilde{w}_{a}^{\ast}(\tilde{p}^{\dagger}))\geqslant 0.$ (53)
Further, from the first-order conditions of optimality for the social cost function, we know that $\tilde{x}^{\dagger}$ is socially optimal if and only if, for every $\tilde{x}\in\tilde{X},$
@@
Using Lemma D.3 in Appendix D, we can equivalently write (54) in terms of edge flows as follows
@@
where $w^{\dagger}$ is the edge flow corresponding to the route flow $x^{\dagger}$. Comparing (53) with (55), we note that $\tilde{w}^{\ast}(p^{\dagger})$ is the minimizer of social cost function $\tilde{\Phi}$. Therefore, $\tilde{w}^{\ast}(\tilde{p}^{\dagger})=\tilde{w}^{\dagger}$.
The proof that $\tilde{\textbf{P}}^{\dagger}$ is a singleton follows by contradiction, which is analogous to that in Proposition 3.1. We omit the details for the sake of brevity.
C.2 Proof of Proposition 4.4
The proof follows by verifying the requirements (R1)-(R2) in Proposition 3.3. Requirement (R2) holds since the strategy space is a compact set. It suffices to show that requirement (R1) holds. Towards this goal, we define a Lyapunov function candidate $V(\tilde{p})=(\tilde{p}-\tilde{p}^{\dagger})^{\top}\Delta(\tilde{p}-\tilde{p}^{\dagger})$ for the dynamical system (38), where $\Delta\in\mathbb{R}^{|\tilde{\mathcal{E}}|\times|\tilde{\mathcal{E}}|}$ is a diagonal matrix defined in (37). Due to the strict monotonicity and convexity of $l_{a}(\cdot)$, it follows that $\Delta_{a,a}>0$ for every $a\in\tilde{\mathcal{E}}$. Consequently, the Lyapunov function candidate is positive definite.
We show that there exists a positive scalar $r$ such that for any $\tilde{p}\in\mathcal{B}_{r}(\tilde{p}^{\dagger})$, the following holds:
@@
$=-2V(\tilde{p})+2\sum_{a\in\tilde{\mathcal{E}}}\Delta_{a,a}(\tilde{p}_{a}-\tilde{p}^{\dagger}_{a})\left(\phi_{a}(\tilde{p})-\phi_{a}(\tilde{p}^{\dagger})\right),$
@@
where for every $a\in\tilde{\mathcal{E}}$, $\phi_{a}(\tilde{p}):=\tilde{w}^{\ast}_{a}(\tilde{p})\nabla l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}))$. Thus, to show local convergence, it suffices to show that there exists $r>0$ such that
@@
$\sum_{a\in\tilde{\mathcal{E}}}\Delta_{a,a}(\tilde{p}_{a}-\tilde{p}^{\dagger}_{a})\left(\phi_{a}(\tilde{p})-\phi_{a}(\tilde{p}^{\dagger})\right)\leqslant 0,\quad\forall\tilde{p}\in\mathcal{B}_{r}(\tilde{p}^{\dagger}).$